# DONOD: Robust and Generalizable Instruction Fine-Tuning for LLMs via Model-Intrinsic Dataset Pruning

### v1

@@ Abstract
Ad-hoc instruction fine-tuning of large language models (LLMs) is widely adopted for domain-specific adaptation. While domain-specific supervised fine-tuning (SFT) is effective and efficient, it often weakens cross-domain generalization and struggles with noisy training data. To address these challenges, we propose DONOD, a lightweight model-intrinsic data pruning method. Our approach evaluates data using two model-parameter-based metrics: Delta of Norm (DON), which captures the cumulative influence on model weights, and Norm of Delta (NOD), which quantifies weight instability. Moreover, by employing the Technique for Order of Preference by Similarity to Ideal Solution (TOPSIS) algorithm, we effectively filter noisy, unlearnable, and generalization-harming samples without relying on auxiliary models during the SFT process. Experiments on mathematical tasks demonstrate that data selected by DONOD achieve superior fine-tuning efficiency and improved robustness against noisy data. By filtering out 70% of the full dataset, we improve target-domain accuracy by 14.90% and cross-domain accuracy by 5.67%. Meanwhile, our selected data present superior cross-architecture generalization. Data pruned by smaller models (e.g., Llama 3.1-8B) generalize effectively on larger models (e.g., Llama 2-13B). Compared to existing related methodologies, DONOD demonstrates comparable or superior performance while remaining dataset-agnostic, enabling broader applicability.
@@ 1 Introduction
Figure 1: Overview performance of our proposed DONOD. Result obtained on the same setting used in § 4.3
Large language models (LLMs) have revolutionized natural language processing, but their effective adaptation to domain-specific tasks remains challenging. Supervised fine-tuning (SFT) is a widely adopted approach for tailoring LLMs to specialized applications. However, it faces critical limitations: overfitting to narrow domains, degradation in cross-domain performance, and sensitivity to noisy or low-quality training data. These issues stem from the inherent trade-off between domain-specific adaptation and generalization, exacerbated by the reliance of existing methods on auxilia-
ry models or predefined validation sets for data selection.
Traditional ad-hoc SFT methods often prioritize domain-specific performance, leading to catastrophic forgetting, where the model loses capabilities in non-target domains. While efforts to mitigate this include dataset curation or mixing diverse data, these approaches either incur high costs or fail to address noise-induced instability. Recent advances in data pruning, such as reward model-guided selection or gradient-based methods, introduce computational overhead or require access to target validation sets, limiting their practicality in real-world scenarios.
To address the issues above, we propose DONOD, a model-intrinsic dataset pruning framework that jointly optimizes for learning efficacy and generalization preservation. DONOD leverages two complementary metrics derived from the model’s weight dynamics. Delta of Norm (DON) captures the cumulative shift in model weight magnitudes during training, acting as a proxy for generalization by favoring samples that stabilize parameter norms. Norm of Delta (NOD): Quantifies the instability caused by individual samples, filtering out noisy or unlearnable data by identifying abrupt weight changes. These metrics are integrated via the Technique for Order of Preference by Similarity to Ideal Solution (TOPSIS) algorithm, which systematically balances the competing goals of maximizing generalization via DON and minimizing harmful updates via NOD. Crucially, DONOD operates without auxiliary models or validation sets, relying solely on the model’s intrinsic training dynamics, thereby ensuring high efficiency.
Experimental evaluations on mathematical reasoning and logical reasoning tasks demonstrate DONOD’s efficacy: pruning 70% of the full dataset improves target-domain accuracy by 14.90% and cross-domain accuracy by 5.67% compared to standard SFT. Notably, datasets pruned by smaller models (e.g., Llama 3.1-8B) generalize effectively to larger architectures (e.g., Llama 2-13B), showcasing strong cross-architecture robustness. DONOD outperforms existing state-of-the-art methods, achieving superior performance with fewer samples while maintaining dataset-agnostic flexibility.
We summarize our contribution as follows: (1) We propose a novel framework for model-intrinsic dataset pruning that unifies generalization and noise filtering through weight-norm analysis. (2) We design a scalable implementation focusing on output-layer weight dynamics, reducing computational overhead while preserving signal fidelity. (3) Empirical validation of cross-domain and cross-architecture generalization demonstrates DONOD’s broad applicability in real-world SFT scenarios.
@@ 2 Related Work
Dataset pruning for LLM SFT critically impacts model performance. Traditional methods often rely on external models as quality judges (Du et al. (2023), Chen et al. (2024)) or employ reward models to identify high-quality data. However, this dependence on auxiliary models—whether trained from scratch or repurposed—incurs significant computational costs and limits scalability. Recent studies (Li et al. (2025)) further question the effectiveness of this paradigm.
@@
Alternative approaches focus on intrinsic data metrics. For instance, Cao et al. (2024) proposes evaluating data quality through features like length, naturalness, and coherence. However, the field lacks consensus on universal metrics: while Chen et al. (2023) emphasizes diversity, Liu et al. (2024) argues for prioritizing complex or challenging samples. This ambiguity motivates the third category—model-intrinsic methods—which leverage the model’s own training dynamics to bypass explicit metric definitions. As noted by Jiang et al. (2019), the model’s response to data inherently signals its utility for learning, enabling automated dataset pruning.
@@
Model-intrinsic pruning methods bifurcate based on scenarios. The first category assumes access to a target data distribution, often via validation or development sets. For example, Mindermann et al. (2022) approximates loss differences between holdout and training sets, while Xia et al. (2024) uses gradient similarity between validation and training data. These methods falter when target distributions are ambiguous or undefined—common in scenarios aiming to enhance broad capabilities rather than optimize for specific benchmarks.
@@
The second category eliminates reliance on target distributions. Works like Wang et al. (2024), Jiang et al. (2019), Loshchilov & Hutter (2016), and Li et al. (2024) employ loss or perplexity thresholds, assuming high-loss samples are valuable learning challenges. However, this assumption proves brittle for noisy or mislabeled data, where high loss reflects annotation errors rather than learnable patterns. Furthermore, excessively difficult samples may exceed the model’s current capacity, rendering them unproductive for training.
@@
Both categories neglect cross-domain generalization. Methods targeting specific distributions risk catastrophic forgetting, where performance gains on target tasks degrade generalizability. Conversely, loss-based selection exacerbates this by prioritizing samples that induce large weight updates, destabilizing pre-trained knowledge.
@@
To address these limitations, we propose DONOD. DON functions as a proxy of generalization, while NOD recognizes that the sample causes significant instability in the model weight. Integrated via the TOPSIS, DONOD filter noisy, unlearnable, and generalization-harming samples without auxiliary models or predefined targets.
3 Methodology
@@ 3.1 Overview
DONOD begins with a critical observation regarding ad-hoc instruction fine-tuning (SFT) for large language models (LLMs). Traditional SFT methods are highly data-dependent, and using a narrowly focused dataset can lead to catastrophic forgetting, which leads to performance degradation in non-target domains during specialization. For instance, fine-tuning on mathematical tasks may significantly impair the model’s ability in reading comprehension. While mixing diverse data can mitigate this issue, it introduces additional training costs and issues of domain proportion balance. Furthermore, the quality of individual samples significantly impacts fine-tuning efficacy, yet identifying high-quality data remains an open problem.
To address these issues, DONOD introduces a data-efficient alternative: instead of augmenting datasets, we prune harmful or low-quality samples that impede generalization. Our approach is grounded in a key insight: the Frobenius norm of weight changes (before vs. after fine-tuning) can estimate each sample’s impact. Specifically, DONOD consists of three core components: 1). DON and NOD. These metrics quantify a sample’s tendency to exacerbate overfitting or catastrophic forgetting. 2). TOPSIS Pruning: Samples are ranked and selected based on DON and NOD, retaining only high-quality data that balances task-specific improvement and cross-domain generalization. 3). Proxy-Based Scalability: We focus on output-layer weight changes as a proxy for full-model behavior, significantly reducing overhead without sacrificing efficacy.
Figure 2: Overview of our proposed DONNOD, which follows a lightweight, three-step pipeline: (1) Compute DON and NOD metrics for each sample, (2) Apply TOPSIS pruning to filter harmful/low-quality data, and (3) Train the model on the pruned dataset.
3.2 Theoretical Foundations
3.2.1 Choice of Frobenius Norm
3.2.2 DON and NOD Metrics
@@
Let $D$ denote an ad-hoc dataset for instruction fine-tuning, for a specific data sample $D_{i}\in D$, let $\{W^{l}\}_{l}^{L}$ represent the weight of the model before fine-tuning and $\{W^{\prime l}\}_{l}^{L}$ the weight matrix after fine-tuning on $D_{i}$. We employ the $DON_{total}$ and $NOD_{total}$ to quantify this change. Specifically:
$\text{DON}_{total}=\sum_{l=1}^{L}(\|W^{l}\|_{F}-\|W^{\prime l}\|_{F})=\sum_{l=1}^{L}(\sqrt{\sum_{i=1}^{m_{l}}\sum_{j=1}^{n_{l}}|w^{l}_{i,j}|^{2}}-\sqrt{\sum_{i=1}^{m_{l}}\sum_{j=1}^{n_{l}}|w’^{l}_{i,j}|^{2}}),$ (1)
where $m_{l}$ and $n_{l}$ are the dimensions of the weight matrix $mW^{l}$ of layer l. Meanwhile, the NOD is defined as:
$\text{NOD}_{total}=\sum_{l=1}^{L}\|W^{l}-W’^{l}\|_{F}=\sum_{l=1}^{L}\sqrt{\sum_{i=1}^{m_{l}}\sum_{j=1}^{n_{l}}\|w^{l}_{i,j}-w’^{l}_{i,j}\|^{2}}.$ (2)
@@
$DON_{total}$ captures the cumulative shift in the model’s weight magnitude, while $NOD_{total}$ measures the direct geometric displacement of weights in parameter space at the current step. These metrics are complementary: $DON_{total}$ reflects the overall scaling of weights, whereas $NOD_{total}$ reflects the sensitivity of model weight on a single sample.
3.2.3 Intuition Behind DON and NOD Metrics
DON as a Proxy for Generalization: A negative DON indicates that fine-tuning on $D_{i}$ increases the model’s weight magnitude. Works from Bartlett (1996) and Yin et al. (2020) show that under certain constraints, smaller norms of weight enable better generalization from the perspective of statistical learning theory. This is also supported by the study of Shalev-Shwartz & Ben-David (2014), its increment relates to the regularization principles, where smaller norms often correlate with lower generalization error (e.g., weight decay). Intuitively, samples that induce a significant increase in the Frobenius norm contribute to the complexity of the model, potentially damage its ability to generalize across domains, and high and positive DON indicate the simplification of the model and better generalization. Our experimental results support this intuition, showing that samples yielding high DON positive values improve cross-domain accuracy.
NOD as an Indicator of Bad Sample: A high NOD value indicates that the data point $D_{i}$ has a significant influence on the model. In the context of ad-hoc SFT, training begins with a model that already possesses a certain level of generalization. Since the model is unlikely to encounter entirely new information or learn fundamentally new concepts after pretraining, data points with high NOD values are often indicative of low-quality samples, such as mislabeled or noisy data. To address this, our method focuses on filtering out samples with high NOD values, thereby removing noisy or unlearnable data from the training process.
3.3 Implementation Details
3.3.1 Focus on Output Layer Weights
@@
While the Frobenius norm can be applied to all model weights, we restrict the use of the weight to the output layer. Prior works (Nadipalli (2025) and Rosati et al. (2024)) have shown that fine-tuning primarily affects later layers, with the output layer acting as a bottleneck for domain adaptation. This restriction benefits in two critical ways: 1). Computational Efficiency: The output layer is typically smaller than the hidden layers, reducing the computational cost of computing norms across iterations, 2). Interpretability: Output layer updates correlate more directly with task performance, avoiding the entangled representations of deeper layers.
By leveraging the output layer’s sensitivity and computational tractability, our approach balances theoretical rigor with practical feasibility, enabling lightweight yet effective data pruning during SFT. Our experiments further validate this choice in § 4.3, showing that data selected based on output layer metrics generalizes effectively across architectures, with pruning decisions from smaller models transferring to larger ones.
Applying this, we obtained a simplified version of (1) and (2) as:
$\text{DON}=\|W^{L}\|_{F}-\|W^{\prime L}\|_{F}=\sqrt{\sum_{i=1}^{m_{L}}\sum_{j=1}^{n_{L}}|w^{L}_{i,j}|^{2}}-\sqrt{\sum_{i=1}^{m_{L}}\sum_{j=1}^{n_{L}}|w^{\prime L}_{i,j}|^{2}},$ (3)
@@
3.3.2 Integration of TOPSIS
TOPSIS is a multi-criteria decision analysis (MCDA) method that ranks alternatives by their relative closeness to an ideal solution. In DONOD, TOPSIS is employed to resolve the inherent tension between the two metrics—DON and NOD—by identifying samples that simultaneously maximize DON (to enhance generalization) and minimize NOD (to avoid noisy or harmful updates). After computing the DON and NOD for each sample, we integrate TOPSIS to rank samples. Details of the algorithm can be found in Appendix  A.5.
TOPSIS inherently balances conflicting objectives (maximizing DON and minimizing NOD) by leveraging geometric distance in the normalized metric space. Moreover, normalization mitigates the impact of differing metric magnitudes, ensuring neither DON nor NOD dominates the ranking. It avoids subjective weight assignment (unlike weighted sum) and provides a total ordering of samples (unlike Pareto optimality), which is critical for deterministic pruning decisions. These factors make TOPSIS considered as particularly suitable here.

### v2

# DONOD: Efficient and Generalizable Instruction Fine-Tuning for LLMs via Model-Intrinsic Dataset Pruning

@@ Abstract
Ad-hoc instruction fine-tuning of large language models (LLMs) is widely adopted for domain-specific adaptation. While domain-specific supervised fine-tuning (SFT) is effective and efficient, it often weakens cross-domain generalization and struggles with noisy training data. To address these challenges, we propose DONOD, a lightweight model-intrinsic data pruning method. Our approach evaluates data using two model-parameter-based metrics: Delta of Norm (DON), which captures the cumulative influence on model weights, and Norm of Delta (NOD), which quantifies weight instability. Moreover, by employing the Technique for Order of Preference by Similarity to Ideal Solution (TOPSIS) algorithm, we effectively filter noisy, unlearnable, and generalization-harming samples without relying on auxiliary models during the SFT process. Experiments on mathematical tasks demonstrate that data selected by DONOD achieves superior fine-tuning efficiency and improved robustness against noisy data. By filtering out 70% of the whole dataset, we improve target-domain accuracy by 14.90% and cross-domain accuracy by 5.67%. Meanwhile, our selected data present superior cross-architecture generalization. Data pruned by smaller models (e.g., Llama 3.1-8B) generalize effectively on larger models (e.g., Llama 2-13B). Compared to existing related methodologies, DONOD demonstrates comparable or superior performance while remaining dataset-agnostic, enabling broader applicability. Code will be made publicly available.
@@ 1 Introduction
In recent years, large language models (LLMs) have demonstrated strong generalization capabilities and remarkable success across a wide range of applications (Achiam et al. 2023; Meta AI 2024; Li et al. 2024b). To further align LLMs with human intent and diverse user needs, supervised fine-tuning (SFT) on curated instruction–response pairs has become a widely adopted approach (Cao et al. 2023; Peng et al. 2023). However, fine-tuning LLMs with massive instruction datasets incurs substantial computational costs. More importantly, recent studies have pointed out that the quality of the fine-tuning data outweighs its quantity (Li et al. 2025b; Xia et al. 2024; Wang et al. 2023). In practice, large-scale instruction data are typically collected via web scraping or weak supervision, which inevitably introduces noisy or redundant samples, thereby hindering the capability (Li and Zhang 2021; Szep et al. 2024; Yang et al. 2024). While manually constructing high-quality SFT datasets is ideal, it is prohibitively expensive at scale. As a practical alternative, recent research focuses on data selection methods that identify high-quality subsets from existing SFT data (Li et al. 2025b; Xia et al. 2024; Wang et al. 2023). Such approaches have shown that models trained on selected subsets can match or even surpass the performance of those trained on the full dataset, offering a promising way to maintain model quality while significantly reducing computational costs.
Recent advances, such as reward-model-based filtering (Xu et al. 2025) and gradient-based pruning (Xia et al. 2024), have pushed this direction forward. Although achieving promising results, many of these methods incur huge computational overhead (Xie et al. 2023) or rely on task-specific validation sets (Xia et al. 2024), which may limit their scalability across diverse domains. Furthermore, recent studies have found that models fine-tuned on such data are prone to domain overfitting, along with degraded generalization across domains (Li and Zhang 2021; Szep et al. 2024), which poses a unique challenge for ad-hoc instruction fine-tuning beyond the general SFT. Therefore, these issues underscore the pressing need for a more principled, data-efficient view to scale LLM fine-tuning more cost-effectively and robustly, with the potential to reduce compute overhead and amplify impact in real-world deployments substantially. This raises a central question: How can we select the most representative and generalizable samples from large-scale instruction datasets to enable efficient and robust LLM fine-tuning?
Figure 1: Performance of our proposed DONOD across various evaluation benchmarks, consistently outperforming other baselines, e.g., all data, despite using only 20% of the training data. These results highlight the potential of principled data pruning for efficient and robust LLM fine-tuning.
To address the issues above, we propose DONOD, a model-intrinsic dataset pruning method that selects a significantly reduced subset of the most representative samples. Specifically, DONOD introduces two complementary metrics derived from the model’s training dynamics: Delta of Norm (DON) and Norm of Delta (NOD). The DON metric captures the cumulative shift in parameter magnitudes throughout training, serving as a proxy for generalization, which encourages the selection of samples that lead to stable norm evolution and contribute to smooth convergence. Meanwhile, NOD measures the instability induced by individual samples by quantifying sample-specific weight fluctuations, thereby identifying noisy or unlearnable data. To reconcile these dual objectives, maximizing generalization via DON and minimizing ambiguous updates via NOD, we employ the Technique for Order of Preference by Similarity to the Ideal Solution (TOPSIS) algorithm (Hwang and Yoon 1981; Chakraborty 2022), which allows principled ranking of samples based on proximity to the ideal pruning criterion. Importantly, DONOD requires no auxiliary models, domain-specific heuristics, or validation sets, relying purely on intrinsic training signals, thus ensuring computational efficiency and broad applicability. As a result, DONOD enables efficient instruction tuning while preserving or even enhancing generalization performance with significantly fewer training examples.
Experiment results across diverse downstream instruction-tuning benchmarks and LLM architectures demonstrate that DONOD consistently outperforms state-of-the-art baselines, achieving superior performance with substantially fewer samples while maintaining dataset-agnostic flexibility. Notably, compared with the full-data SFT setting, DONOD improves target-domain accuracy by 14.90% and cross-domain accuracy by 5.67% when using only 30% of the data. Moreover, we demonstrate strong cross-architecture generalization, where data selected by lighter-weight models (e.g., LLaMA 3.1-8B (Grattafiori et al. 2024)) can generalize well across larger or different architectures, such as LLaMA 2-13B (Touvron et al. 2023) and Qwen 2.5-7B (Team 2024), etc. Additionally, we further validate the robustness of DONOD in more complex and realistic noisy settings, highlighting its practical significance. We summarize our contribution as follows:
  • We propose DONOD, a lightweight, model-intrinsic data selection framework for accelerating the fine-tuning of LLMs.
  • We introduce two complementary metrics, DON and NOD, where DON captures the generalization impact of samples and NOD identifies noisy or unlearnable samples.
  • Experimental results validate that our method outperforms existing methods in terms of cross-domain and cross-architecture generalization performance, e.g., achieving lossless training acceleration with only 20% of the data, highlighting the scalability and robustness of our approach.
@@ 2 Related Work
Dataset pruning for supervised fine-tuning on LLMs critically impacts model performance. Traditional methods often rely on external models as quality judges ((Du, Zong, and Zhang 2023), (Chen et al. 2024)) or employ reward models to identify high-quality data (Yang et al. 2025a). However, this dependence on auxiliary models, whether trained from scratch or repurposed, incurs significant computational costs and limits scalability. Recent studies ((Li et al. 2025a)) further question the effectiveness of this paradigm.
@@
Alternative approaches focus on intrinsic data metrics. For instance, (Cao et al. 2024) proposes evaluating data quality through features like length, naturalness, and coherence. However, the field lacks consensus on universal metrics: while (Chen et al. 2023) emphasizes diversity, (Liu et al. 2024) argues for prioritizing complex or challenging samples. This ambiguity motivates the third category—model-intrinsic methods—which leverage the model’s training dynamics to bypass explicit metric definitions. As noted by (Jiang et al. 2019; Yang et al. 2025b), the model’s response to data inherently signals its utility for learning, enabling automated dataset pruning.
Figure 2: Overview of our proposed DONOD, which follows a lightweight, three-step pipeline: (1) Compute DON and NOD metrics for each sample, (2) Apply TOPSIS pruning to select representative data and filter harmful/low-quality data, and (3) Train the model on the pruned dataset.
@@
Model-intrinsic pruning methods branch out based on various scenarios. The first category assumes access to a target data distribution, often via validation or development sets. For example, (Mindermann et al. 2022) approximates loss differences between holdout and training sets, while (Xia et al. 2024) uses gradient similarity between validation and training data. These methods falter when target distributions are ambiguous or undefined, which is common in scenarios aiming to enhance broad capabilities rather than optimize for specific benchmarks.
@@
The second category eliminates reliance on target distributions. Works like (Wang et al. 2024), (Jiang et al. 2019), (Loshchilov and Hutter 2016), and (Li et al. 2024b) employ loss or perplexity thresholds, assuming high-loss samples are valuable learning challenges. However, this assumption proves brittle for noisy or mislabeled data (Yang et al. 2024), where high loss reflects annotation errors rather than learnable patterns. Furthermore, challenging samples may exceed the model’s current capacity, rendering them unproductive for training.
@@
Both categories neglect cross-domain generalization. Methods targeting specific distributions risk catastrophic forgetting, where performance gains on target tasks degrade generalizability. Conversely, loss-based selection exacerbates this by prioritizing samples that induce significant weight updates, destabilizing pre-trained knowledge.
@@
To address these limitations, we propose DONOD. DON functions as a proxy of generalization, while NOD recognizes that the sample causes significant instability in the model weight. Integrated via the TOPSIS, DONOD filters noisy, unlearnable, and generalization-harming samples without auxiliary models or predefined targets.
3 The Proposed Method
@@ 3.1 Overview
Traditional SFT methods are highly data-dependent, and using a narrowly focused dataset can lead to catastrophic forgetting, resulting in performance degradation in non-target domains during specialization. For instance, fine-tuning on mathematical tasks may significantly impair the model’s ability in reading comprehension. While mixing diverse data can mitigate this, it introduces additional training costs and issues of domain proportion balance. Furthermore, the quality of individual samples significantly impacts fine-tuning efficacy, yet identifying high-quality data remains an open problem.
To this end, DONOD introduces a data-efficient alternative to select a reduced subset of representative samples and achieve SFT acceleration. Specifically, DONOD consists of three core components: 1) DON and NOD metrics based on the Frobenius norm are used to estimate the samples’ impact on model weight update. 2) TOPSIS is a multi-objective decision mechanism that balances task-specific gain and cross-domain generalization, ensuring that the selected subset preserves both in-domain effectiveness and robustness to distribution shifts. 3) By approximating the full model behavior through changes in the output layer, DONOD enables lightweight and scalable selection without the need to backpropagate through the entire model. These components form an efficient data selection framework that supports accurate, low-cost subset selection, enabling fine-tuning with significantly fewer samples while maintaining or even improving model performance.
3.2 DON and NOD Metrics
@@
Let $D$ denote an ad-hoc dataset for instruction fine-tuning. For a specific data sample $D_{i}\in D$, let $\{W^{l}\}_{l}^{L}$ represent the weight of the model before fine-tuning and $\{W^{\prime l}\}_{l}^{L}$ the weight matrix after fine-tuning on $D_{i}$. We employ the DON and NOD to quantify this change. Specifically, the DON is defined as:
DON $=\sum_{l=1}^{L}\left(\|W^{l}\|_{F}-\|W^{\prime l}\|_{F}\right)$ (1)
$=\sum_{l=1}^{L}\left(\sqrt{\sum_{i=1}^{m_{l}}\sum_{j=1}^{n_{l}}|w^{l}_{i,j}|^{2}}-\sqrt{\sum_{i=1}^{m_{l}}\sum_{j=1}^{n_{l}}|w^{\prime l}_{i,j}|^{2}}\right),$
where $m_{l}$ and $n_{l}$ are the dimensions of the weight matrix $W^{l}$ of layer $l$, $\|\cdot\|_{F}$ denotes the Frobenius norm. Here, we adopt the Frobenius norm due to its ability to capture fine-grained structural changes across all weight elements while maintaining computational efficiency, making it a suitable and scalable proxy for quantifying sample-level influence in large-scale models. Thus, DON captures the cumulative shift in the model’s weight magnitude. From a generalization perspective, a positive DON suggests that a sample reduces the model’s Frobenius norm, which is associated with lower complexity and better generalization (Bartlett 1996; Yin, Ramchandran, and Bartlett 2020; Shalev-Shwartz and Ben-David 2014). Meanwhile, the NOD is defined as:
$\text{NOD}=\sum_{l=1}^{L}\|W^{l}-W’^{l}\|_{F}=\sum_{l=1}^{L}\sqrt{\sum_{i=1}^{m_{l}}\sum_{j=1}^{n_{l}}\|w^{l}_{i,j}-w’^{l}_{i,j}\|^{2}},$ (2)
@@
which measures the direct geometric displacement of weights in parameter space at the current step. Thus, in the context of SFT, NOD quantifies how drastically a single sample perturbs the parameter space. These metrics are complementary: DON reflects the overall scaling of weights, whereas NOD reflects the sensitivity of model weight on a single sample.
@@
While the Frobenius norm can be applied to whole model weights, prior works (Nadipalli 2025; Rosati et al. 2024) show that fine-tuning primarily affects later layers, with the output layer acting as a bottleneck for domain adaptation. Therefore, we estimate the sample influence using the weights of the last layer, which brings two benefits: 1) Computational Efficiency: The output layer is typically smaller than the hidden layers, reducing the computational cost of computing norms across iterations, 2) Interpretability: Output layer updates correlate more directly with task performance, avoiding the entangled representations of deeper layers.
In this way, we derive a simplified version of the computation of (1) and (2):
$\begin{split}\text{DON}&=\|W^{L}\|_{F}-\|W^{\prime L}\|_{F}\\ &=\sqrt{\sum_{i=1}^{m_{L}}\sum_{j=1}^{n_{L}}|w^{L}_{i,j}|^{2}}-\sqrt{\sum_{i=1}^{m_{L}}\sum_{j=1}^{n_{L}}|w^{\prime L}_{i,j}|^{2}},\end{split}$ (3)
@@
3.3 Integration of TOPSIS
TOPSIS is a multi-criteria decision analysis (MCDA) method that ranks alternatives by their relative closeness to an ideal solution. In DONOD, TOPSIS is employed to resolve the inherent tension between the two metrics—DON and NOD—by identifying samples that simultaneously maximize DON (to enhance generalization) and minimize NOD (to avoid noise). After computing the DON and NOD for each sample, we integrate TOPSIS to rank the data points.
TOPSIS inherently balances conflicting objectives, i.e., maximizing DON and minimizing NOD, by leveraging geometric distance in the normalized metric space. Moreover, normalization mitigates the impact of differing metric magnitudes, ensuring neither DON nor NOD dominates the ranking. It avoids subjective weight assignment (unlike weighted sum) and provides a total ordering of samples (unlike Pareto optimality), which is critical for deterministic pruning decisions. Thus, we choose TOPSIS in our framework.
Computational Complexity
The algorithm’s computational complexity consists of: (1) per-sample DON and NOD computation and (2) TOPSIS-based sample selection.
In the first phase, we perform a forward pass ($O(P)$ time per sample, where $P$ is the model’s parameter count), a backward pass restricted to the output layer $(O$)O(O) time, $O$ being the output layer’s parameters), and compute Frobenius norms for weight updates $(O$)O(O)). Since $O\subset P$, the per-sample cost simplifies to $O(P)$, resulting in a total training complexity of $O(N\cdot P)$, where $N$ is the number of training samples.